A person invests $8000 for one year. Part of this amount is invested at 2.65% simple interest and the rest is invested at 3.12% simple interest. If the total interest received from both investments is $235.50, find the amounts invested at each rate
Answer by stanbon(75887)

A person invests $8000 for one year. Part of this amount is invested at 2.65% simple interest and the rest is invested at 3.12% simple interest. If the total interest received from both investments is $235.50, find the amounts invested at each rate.
------------------
Equations:
Quantity Equation: x + y = 8000
Interest Equation:0.0265x + 0.0312y = 235.50
-------------------------
Multiply thru the 1st equation by 265
Multiply thru the 2nd equation by 10000
265x + 265y = 265*8000
265x + 312y = 2355000
----------------
Subtract 1st Eq. from 2nd and solve for "y":
47y = 235000
y = $5000 (amt. invested at 2.65%)
---
x = 8000-5000 = $3000 (amt. invested at 3.12%)
==========================
Cheers,
Stan H.
===============